title="Colnbrook with Poyle Parish Council in Berkshire">

Biffa Environmental GroupTue, 21st November 2017

Biffa Environmental Group

 

BIFFA SITE MEETING

Thursday 15th March 2012

 

It was confirmed that waste operations ceased in September 2011, and that all transport movement since have been for the importation of top soil to be used for the restoration programme.

 

March/April will see the commencement of the final phases of the capping off programme which will see some 70,000 cubic meters of capping material laid, and between 60 and 80 additional gas wells installed. The gas is converted on site into electricity and fed into the National Grid. This aspect of the closing proceedure should be completed by late June early July.

 

Once this "capping off" process is completed the restoration of the land to agricultural grazing land will begin. This involves the laying of uncontaminated top soil over the topping off material and this process should be completed by the end of September / early October 2012.

 

Biffa Landfill Liaison Meeting – 14th July 2011

Operational Matters.

Andy White stated that capping and drilling of cell 13 had commenced. Tipping into cell 12, the area that the Biogenie occupied, is continuing.

The upgrading of the gas pipes had caused some odour problems whilst the system was opened to add ‘Tees’ into the system. Some complaints were received both direct and via the E.A. explanations were given and the problem was quickly resolved.

The leachate treatment plant is now up and running and is now covered by monitored CCTV to prevent cable theft. A revised leachate management plan has been submitted to the E.A.

Although reported at the last meeting waste levels had fallen to all-time lows, it is now reported that they have picked up to approximately 1100 tonne per day. This will see the site fulfilling its commitment of completion by December 2011. It will be observed shortly that less of the waste material vehicles will be entering the site and more of the smaller rigid tippers bringing in restoration materials. Most of the restoration material, currently being stockpiled, is being sourced from T2 & Kensington.

Chris Lowe (E.A.) stated that land to the north of the site was heavily contaminated not within control of the Biffa site, but from an old working. He also stated that the land which Goodman’s wish to use may have to be separately licenced if they win their application, and proceed to level the land for the railway infrastructure. The old landfill workings will be exposed in levelling, and will then be classified as current landfill. The only alternative would be to ‘level up’ which may become impractical.

Site inspection meeting by both SBC and the E.A. are to be scheduled coincidental in order to reduce the time site management is involved in the process.

Roger Kirkham (SBC Planning) highlighted some implications that the Goodman’s application will impact upon the site. Bore holes to take samples from the south west corner of the boundary to the land which is within Goodman’s control.

SBC is pulling out of the Berkshire Minerals & Waste Policy Strategy.

The next schedule meeting would have been in November, but due to the scheduled site completion it was brought forward to be 20th October @ 10.00am.